Milwaukee Tool has the following​ stockholders' equity account. The​ firm's common stock currently sells for ​$3.78 per share.

Preferred stock                                                                 $93,000

Common stock (400,000 shares at $1.01 par)                   404,000

Paid-in capital in excess of par                                          215,000

Retained earnings                                                             320,000

Total stockholders' equity                                                  1,032,000

a.  The balance in preferred stock after the ​$0.01 cash dividend is $_______ (Round to the nearest​ dollar.)

The balance in common stock after the $0.01cash dividend is ​$________ (Round to the nearest​ dollar.)

The balance in​ paid-in capital after the $0.01cash dividend is $________ ​(Round to the nearest​ dollar.)

The balance in retained earnings after the ​$0.01 cash dividend is $_______ (Round to the nearest​ dollar.)

The balance in total​ stockholders' equity after the $0.01 cash dividend is ​$__________ (Round to the nearest​ dollar.)

b.  The balance in preferred stock after the 20% stock dividend is $_________ (Round to the nearest​ dollar.)

The balance in common stock after the 20​% stock dividend is ​$__________ (Round to the nearest​ dollar.)

The balance in​ paid-in capital after the 20% stock dividend is $__________​(Round to the nearest​ dollar.)

The balance in retained earnings after the 20% stock dividend is $_________ (Round to the nearest​ dollar.)

The total​ stockholder's equity after the 20% stock dividend is ​$___________(Round to the nearest​ dollar.)
Answer #1

Answer 1:

Cash dividend = $0.01 per share

Hence total dividend paid = $.0.01 * 400000 = $4,000

Retained earnings will reduce by $4,000 and balance in retained earnings = 320,000 - 4000 = $316,000

Preferred stock, Common stock and Paid-in capital in excess of par balances will remain same

Balance in total​ stockholders' equity will be = 93000 + 404000 + 215,000 + 316000 = $1,028,000

As such:

The balance in preferred stock after the ​$0.01 cash dividend is $93,000

The balance in common stock after the $0.01 cash dividend is​ $404,000

The balance in​ paid-in capital after the $0.01 cash dividend is $215,000

The balance in retained earnings after the ​$0.01 cash dividend is $316,000

The balance in total​ stockholders' equity after the $0.01 cash dividend is ​$1,028,000

Answer b:

On 20% stock dividend, new shares issued = 20% * 400000 = 80,000

Since it is small stock dividend:

Retained earnings will reduce by = new shares issued * market price = 80000 * 3.78 = $302,400

New balance in retained earnings = 320000 -302400 = $17,600

Common stock will increase by = 80000 * $1.01 = $80,800

Balance in common stock will be =404000 + 80800 =$484,800

Paid-in capital in excess of par will increase by = 80000 * (3.78 - 1.01) = $221,600

Balance in Paid-in capital in excess of par will be = 215000 + 221600 = $436,600

Balance in total​ stockholders' equity will be = 93000 + 484800 + 436,600 + 17600 = $1,032,000

Hence:

The balance in preferred stock after the 20% stock dividend is $93,000

The balance in common stock after the 20​% stock dividend is ​$484,800

The balance in​ paid-in capital after the 20% stock dividend is $436,600

The balance in retained earnings after the 20% stock dividend is $17,600

The total​ stockholder's equity after the 20% stock dividend is​ $1,032,000